### v1

# Optimal Neural Network Approximation for High-Dimensional Continuous Functions

Abstract.
Recently, the authors of [23] developed a neural network with width $36d(2d+1)$ and depth $11$, which utilizes a special activation function called the elementary universal activation function, to achieve the super approximation property for functions in $C([a,b]^{d})$. That is, the constructed network only requires a fixed number of neurons to approximate a $d$-variate continuous function on a $d$-dimensional hypercube with arbitrary accuracy. Their network uses $\mathcal{O}(d^{2})$ fixed neurons. One natural question to address is whether we can reduce the number of these neurons in such a network. By leveraging a variant of the Kolmogorov Superposition Theorem, our analysis shows that there is a neural network generated by the elementary universal activation function with only $365366d+$ 365 fixed, intrinsic (non-repeated) neurons that attains this super approximation property. Furthermore, we present a family of continuous functions that requires at least width $d$, and therefore at least $d$ intrinsic neurons, to achieve arbitrary accuracy in its approximation. This shows that the requirement of $\mathcal{O}(d)$ intrinsic neurons is optimal in the sense that it grows linearly with the input dimension $d$, unlike some approximation methods where parameters may grow exponentially with $d$.
Key words and phrases:
Neural network, approximation theory, high-dimensional functions, special activation functions
2020 Mathematics Subject Classification:
41A99, 68T07
1. Introduction
The wide applicability of neural networks has generated tremendous interest, leading to many studies on their approximation properties. Some early work on this subject can be traced back to [3, 7]. As summarized in [23], there have been several research paths in this area such as finding nearly optimal asymptotic approximation errors of ReLU networks for various classes of functions [6, 20, 25, 27], deriving nearly optimal non-asymptotic approximation errors for continuous and $C^{s}$ functions [16, 21], mitigating the curse of dimensionality in certain function spaces [1, 4, 18], and improving the approximation properties by using a combination of activation functions and/or constructing more sophisticated ones [17, 22, 23, 26, 27].
Building on the last point, [26] presented several explicit examples of superexpressive activation functions, which if used in a network allows us to approximate a $d$-variate continuous function with a fixed architecture and arbitrary accuracy. That is, the number of neurons remains the same, but the values of the parameters may change. This approach is notably different from using a standard network with commonly used activation functions such as ReLU. To achieve the desired accuracy, a standard network with a commonly used activation function typically needs to have its width and/or its depth increased based on the target accuracy. The growth of the number of neurons in terms of the target accuracy may range from polynomial to, in the worst-case scenario, exponential.
The existence of a special activation function mentioned earlier has been known since the work of [17]; however, its explicit form is unknown, even though the activation itself has many desirable properties such as sigmoidal, strictly increasing, and analytic. In the same vein, [23] introduced several new explicit activation functions, called universal activation functions, that allow a network with a fixed architecture to achieve arbitrary accuracy when approximating a $d$-variate continuous function. In a follow-up work, [24] presented additional examples of universal activation functions and evaluated their performance on various datasets.
More recently, there have also been studies that look into the minimum required width of networks generated by various activation functions to achieve the universal approximation property for functions in $L_{p}$ spaces and continuous functions [2, 5, 8, 10, 14, 15, 19]. We briefly review relevant results for continuous functions. Suppose that $K$ is a compact domain in $\mathbb{R}^{d_{x}}$ and let $w_{\min}$ denote the minimum width. The authors of [5] found that a ReLU network requires $w_{\min}=d_{x}+1$ for functions in $C(K,\mathbb{R})$ (or equivalently, $C(K)$). More generally, for functions in $C(K,\mathbb{R}^{d_{y}})$, a ReLU+STEP network requires $w_{\min}=\max(d_{x}+1,d_{y})$ [19], a network with an arbitrary activation function requires $w_{\min}\geq\max(d_{x},d_{y})$ [2], a ReLU+FLOOR network requires $w_{\min}\geq\max(d_{x},d_{y},2)$ [2], and a Leaky-ReLU network requires $w_{\min}=\max(d_{x}+1,d_{y})+\mathbbm{1}_{d_{y}=d_{x}+1}$ [14]. If we consider functions in $C([0,1]^{d_{x}},\mathbb{R}^{d_{y}})$, then a network generated by an activation function that can be approximated by a sequence of continuous one-to-one continuous functions requires $w_{\min}\geq d_{x}+1$ [8], a network generated by a non-polynomial activation function that is continuously differentiable at least a point requires $w_{\min}\leq d_{x}+d_{y}+1$ [10], and a network with a non-affine polynomial activation function requires $w_{\min}\leq d_{x}+d_{y}+2$ [10].
To approximate a $d$-variate continuous function on a $d$-dimensional hypercube, some network constructions [13, 18, 23] rely on the Kolmogorov Superposition Theorem (KST) [12]. KST represents such a function in terms of compositions and additions of univariate continuous functions on bounded intervals, thereby making the analysis of such a function highly convenient.
The present paper is motivated by the findings of [23]. Their network is capable of approximating functions in $C([a,b]^{d})$ with arbitrary accuracy, using only a width of $36d(2d+1)$ and a depth of $11$, resulting in $\mathcal{O}(d^{2})$ fixed neurons. The authors used the original KST [12] to convert the analysis of a $d$-variate continuous function into that of several univariate continuous functions. Furthermore, they constructed an elementary universal activation function (EUAF) network to approximate a univariate continuous function with arbitrary accuracy. A natural question that arises from their network is whether the same super approximation property can be achieved with only $\mathcal{O}(d)$ fixed neurons, similar to the result obtained by [17]. Although the network constructed in [17] only requires $39d+$ 3 fixed neurons, its drawback is that its activation function is not explicitly known.
The main contributions of this paper are twofold. Firstly, we provide a positive answer to the foregoing question. More specifically, we can show that there is an EUAF network with $365366d+$ 365 intrinsic (non-repeated) neurons that achieves the desired super approximation property. That is, we can approximate a target function in $C([a,b]^{d})$ with arbitrary accuracy using only $\mathcal{O}(d)$ fixed neurons. To obtain a better approximation, only the values of these parameters change. This is in stark contrast to a standard network generated by commonly used activation functions such as ReLU, where the number of parameters typically grows to obtain a more accurate approximation. We have identified that the primary cause of the $\mathcal{O}(d^{2})$ neurons requirement in [23] stems from the original KST needing $12d+$ 1 outer functions and $(2d+1)(d+1)$ inner functions. Using a variant of KST ([9] or [17, Theorem 5]) allows us to only use 1 outer function and $12d+$ 1 inner functions. From the approximation point of view, we just need to approximate these $22d+$ 2 functions by EUAF networks once and evaluate them repeatedly (see Fig. 2). Thus, not only can we bring down the number of intrinsic neurons to $\mathcal{O}(d)$, but the proof is simplified. Secondly, we present a family of continuous functions that requires at least width $d$, and thus at least $d$ intrinsic neurons to achieve arbitrary accuracy in its approximation. Combining these results, we conclude that the requirement of $\mathcal{O}(d)$ fixed intrinsic neurons for approximating functions in $C([a,b]^{d})$ is optimal in the sense that the number of neurons only linearly depends on the input dimension $d$. This requirement is significantly less severe compared to some other approximation methods, which may use an exponentially growing number of parameters. To better understand how our study compares to others in terms of the width and depth requirements, we refer readers to Fig. 1.
The organization of this paper is as follows. In Section 2, we review some basic notations and key ingredients used in the proof of our main results. In Section 3, we show the existence of an EUAF network with $\mathcal{O}(d)$ intrinsic neurons that can approximate functions in $C([a,b]^{d})$ with arbitrary accuracy. Additionally, we present a family of continuous functions that requires at least width $d$ (or $d$ intrinsic neurons) for its network approximation to achieve arbitrary accuracy.
Figure 1. A diagram showing some relevant studies that discuss the requirements for achieving a target approximation accuracy $\epsilon$ in terms of the width, depth, and the input dimension $d$ for a $d$-variate continuous function on a $d$-dimensional hypercube with a scalar output. Studies with width $\mathcal{O}(f_{3}(d,\epsilon))$ and/or depth $\mathcal{O}(f_{4}(d,\epsilon))$ imply that achieving a more accurate approximation requires increasing the depth and/or width based on $d$ and $\epsilon$ (as well as possibly other factors like the modulus of continuity of the target function). The results of our paper and [17, 23, 26] have the super approximation property in the sense that only a fixed number of neurons is needed to achieve arbitrary accuracy. The network constructed in [26] has width $\mathcal{O}(f_{1}(d))$ and depth $\mathcal{O}(f_{2}(d))$, where $f_{1}$ and $f_{2}$ are some functions that are not known explicitly. Our network uses $\mathcal{O}(d)$ fixed intrinsic neurons and an explicitly known activation function (unlike [17]). The width of our network is $\mathcal{O}(d)$ if we view it as in Fig. 2.
2. Preliminaries
For a given activation function $\sigma$, the function $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a $\sigma$ network with $L\in\mathbb{N}$ layers if
$\phi:=\mathbf{\mathcal{L}}_{L}\circ\sigma\circ\mathbf{\mathcal{L}}_{L-1}\circ\sigma\circ\dots\circ\sigma\circ\mathbf{\mathcal{L}}_{1}\circ\sigma\circ\mathbf{\mathcal{L}}_{0},$ (2.1)
where $\mathbf{\mathcal{L}}_{i}(\mathbf{y}):=\mathbf{W}_{i}\mathbf{y}+\mathbf{b}_{i}$ with $\mathbf{W}_{i}\in\mathbb{R}^{N_{i+1}\times N_{i}}$, $\mathbf{y}\in\mathbb{R}^{N_{i}}$, $\mathbf{b}_{i}\in\mathbb{R}^{N_{i+1}}$, $N_{0}=d$, $N_{L+1}=1$, and for a given $\mathbf{y}=(y_{1},\ldots,y_{N_{i}})^{\mathsf{T}}$, the activation function is applied elementwise (i.e., $\sigma(\mathbf{y}):=(\sigma(y_{1}),\ldots,\sigma(y_{N_{i}}))^{\mathsf{T}}$). If $N_{i}=N$ for all $1\leq i\leq L$ (i.e, there are $N$ neurons for each hidden layer), then we say that such a $\sigma$ network has width $N$. Recall that the EUAF introduced in [23] is defined as
$\sigma:=\begin{cases}\left|x-2\lfloor\tfrac{x+1}{2}\rfloor\right|,&x\in[0,\infty),\\ \frac{x}{|x|+1},&x\in(-\infty,0).\\ \end{cases}$ (2.2)
In this paper, we shall focus on EUAF networks (i.e., $\sigma$ networks with $\sigma$ defined in (2.2)).
We first present the two key ingredients in the construction an EUAF network with fixed $\mathcal{O}(d)$ intrinsic (non-repeated) neurons, which can approximate any function in $C([a,b]^{d})$ with arbitrary accuracy. The first ingredient is a version of KST [12], which was studied in [9] and utilized in the construction of the network in [17]. In contrast to the original KST, the following version requires only one outer function and $12d+$ 1 inner functions, which enables us to use only $\mathcal{O}(d)$ fixed, intrinsic neurons.
Theorem 2.1.
([9] or [17, Theorem 5]) Let $\mathbf{x}=(x_{1},\dots,x_{d})^{{}^{\mathsf{T}}}$. There exist $d$ constants $\lambda_{j}>0$, $j=1,\ldots,d$, $\sum_{j=1}^{d}\lambda_{j}\leq 1$, and $12d+$ 1 continuous strictly increasing functions $h_{i}$, $11\leq i\leq 2d+$ 1, which map $[0,1]$ to itself, such that every continuous function $f$ of $d$ variables on $[0,1]^{d}$ can be represented in the form
$f(\mathbf{x})=\sum_{i=1}^{2d+1}g\left(\sum_{j=1}^{d}\lambda_{j}h_{i}(x_{j})\right)$
for some $g\in C([0,1])$ depending on $f$.
The second ingredient is the following result from [23] on the existence of an EUAF network with fixed width and depth that can approximate any function in $C([a,b])$ with arbitrary accuracy.
Theorem 2.2.
[23, Theorem 6] Let $f\in C([a,b])$. Then, for an arbitrary $\epsilon>0$, there exists a function $\phi$ generated by an EUAF network with width 36 and depth 5 such that
$|\phi(x)-f(x)|<\epsilon\quad\text{for any }x\in[a,b]\subseteq\mathbb{R}.$
As outlined in [23], the construction of such an EUAF network was performed by using a three-step procedure: (1) divide the bounded interval into several sub-intervals (the number of these smaller intervals depends on the prescribed error and the target function), (2) build a sub-network that maps each sub-interval to an integer value, and (3) build another sub-network that maps the index of the sub-interval to a function value.
There is also a possibility of reducing the number of neurons in the network, by combining the EUAF with superexpressive activation functions presented in [26]. However, for simplicity, we choose to use the same activation function throughout the entire network and adhere to Theorem 2.2. Moreover, in the context of techniques used in the paper, reducing the number of neurons in the approximation of a function in $C([a,b])$ will not lead to a reduction in the order of magnitude with respect to $d$, when we combine this result with the KST to approximate a function in $C([a,b]^{d})$.
3. Main results
Now, we are ready to present our first main result. The following theorem guarantees the existence of an EUAF network with fixed $\mathcal{O}(d)$ intrinsic (non-repeated) neurons that can approximate any function in $C([a,b]^{d})$ with arbitrary accuracy.
Theorem 3.1.
Let $f\in C([a,b]^{d})$. Then, for an arbitrary $\epsilon>0$, there exists a function $\phi$ generated by an EUAF network with $365366d+$ 365 intrinsic neurons such that
$|f(\mathbf{x})-\phi(\mathbf{x})|<\epsilon\quad\text{for all }\mathbf{x}\in[a,b]^{d}.$
Proof.
Let $\mathbf{x}=(x_{1},\ldots,x_{d})^{\textsf{T}}$ and $\mathbf{y}=(y_{1},\ldots,y_{d})^{\textsf{T}}$. Define ~$\tilde{\mathcal{L}}(t):=(t-a)/(b-a)$ for $t\in[a,b]$ and $\tilde{f}:=f\circ\tilde{\mathcal{L}}$. Clearly, $\tilde{f}\in C([0,1]^{d})$. By Theorem 2.1, we have
$\tilde{f}(\mathbf{y})=\tilde{f}(y_{1},\ldots,y_{d})=\sum_{i=1}^{2d+1}g\left(\sum_{j=1}^{d}\lambda_{j}\tilde{h}_{i}(y_{j})\right),\quad y_{j}\in[0,1]\text{for all }1\leq j\leq d,$ (3.1)
where $\sum_{j=1}^{d}\lambda_{j}\leq 1$ with $\lambda_{j}>0$ for each $j$, each $\tilde{h}_{i}$ is a continuous strictly increasing function mapping $[0,1]$ to itself, and $g\in C([0,1])$. If we define $\mathcal{L}(t):=a+(b-a)t$ for $t\in[0,1]$, then (3.1) yields
$f(x_{1},\ldots,x_{d})=f(\mathcal{L}(y_{1}),\ldots,\mathcal{L}(y_{d}))=\tilde{f}(y_{1},\ldots,y_{d})=\sum_{i=1}^{2d+1}g\left(\sum_{j=1}^{d}\lambda_{j}\tilde{h}_{i}(y_{j})\right)=\sum_{i=1}^{2d+1}g\left(\sum_{j=1}^{d}\lambda_{j}h_{i}(x_{j})\right),$
where $h_{i}:=\tilde{h}_{i}\circ\mathcal{L}$ and $x_{j}\in[a,b]$ for all $i,j$. Note that each $h_{i}$ is now a continuous function mapping $[a,b]$ to $[0,1]$.
Now, arbitrarily fix $\epsilon>0$. First, we focus on the approximation of the outer function $g$. Since $g$ is a uniformly continuous function on $[0,1]$, we know that there exists $\delta>0$ such that
$|g(z_{1})-g(z_{2})|<\frac{\epsilon}{2(2d+1)}\quad\text{for all }z_{1},z_{2}\in[0,1]\text{ with }|z_{1}-z_{2}|<\delta.$ (3.2)
Additionally, by Theorem 2.2, we know that there is an EUAF network $\tilde{\phi}$ with width 36 and depth 5 such that
$|g(z)-\tilde{\phi}(z)|<\frac{\epsilon}{2(2d+1)}\quad\text{for all }z\in[0,1].$ (3.3)
Next, we turn to the approximation of each inner function $h_{i}$. For each $i$, we know by Theorem 2.2 again that there is an EUAF network $\tilde{\psi}_{i}$ with width $36$ and depth $5$ such that
$|h_{i}(z)-\tilde{\psi}_{i}(z)|<\delta\quad\text{for all }z\in[a,b].$ (3.4)
Define $\psi_{i}:=\min\{\max\{\tilde{\psi}_{i},0\},1\}$. If $\tilde{\psi}_{i}(z)<0$ for some $z\in[a,b]$, then
$h_{i}(z)-\psi_{i}(z)<h_{i}(z)-\tilde{\psi}_{i}(z)<\delta,$
since $h_{i}$ is a strictly increasing continuous function whose range is $[0,1]$. Otherwise, if $\tilde{\psi}_{i}(z)>1$ for some $z\in[a,b]$, then
$\psi_{i}(z)-h_{i}(z)<\tilde{\psi}_{i}(z)-h_{i}(z)<\delta$
due to the same reason. Thus, we have
$|h_{i}(z)-\psi_{i}(z)|\leq|h_{i}(z)-\tilde{\psi}_{i}(z)|<\delta\quad\text{for all }z\in[a,b].$
By (2.2), we observe that
$\min\{\max\{t,0\},1\}=\frac{1}{2}\left((t+1)-\sigma(t+1)\right)=\frac{3}{2}\sigma\left(\frac{1}{3}t+\frac{1}{3}\right)-\frac{1}{2}\sigma(t+1)\quad\text{for all }t\in[-1,2],$
which implies that $\psi_{i}$ can be constructed by three more neurons to further process the output of $\tilde{\psi}_{i}$. One neuron is used to compute $\sigma(\tfrac{1}{3}\tilde{\psi}_{i}+\frac{1}{3})$, another is used to compute $\sigma(\tilde{\psi}_{i}+1)$, and the other is used to compute a linear combination of the former two to produce the desired output. Since $\sum_{j=1}^{d}\lambda_{j}=1$ with $\lambda_{j}>0$ and the range of $\psi_{i}$ is $[0,1]$ for all $11\leq i\leq 2d+$ 1, we can immediately see that the range of $\sum_{j=1}^{d}\lambda_{j}\psi_{i}(x_{j})$ is $[0,1]$.
Define
$\phi(\mathbf{x}):=\sum_{i=1}^{2d+1}\tilde{\phi}\left(\sum_{j=1}^{d}\lambda_{j}\psi_{i}(x_{j})\right),\quad\mathbf{x}\in[a,b]^{d}.$
By (3.4), we observe that
$\left|\sum_{j=1}^{d}\lambda_{j}h_{i}(x_{j})-\sum_{j=1}^{d}\lambda_{j}\psi_{i}(x_{j})\right|\leq\sum_{j=1}^{d}\lambda_{j}|h_{i}(x_{j})-\psi_{i}(x_{j})|<\sum_{j=1}^{d}\lambda_{j}\delta=\delta,\quad 1\leq i\leq j.$ (3.5)
Therefore, for $\mathbf{x}\in[a,b]^{d}$, we have
$|f(\mathbf{x})-\phi(\mathbf{x})|=\left|\sum_{i=1}^{2d+1}g\left(\sum_{j=1}^{d}\lambda_{j}h_{i}(x_{j})\right)-\sum_{i=1}^{2d+1}\tilde{\phi}\left(\sum_{j=1}^{d}\lambda_{j}\psi_{i}(x_{j})\right)\right|$
$\leq\left|\sum_{i=1}^{2d+1}g\left(\sum_{j=1}^{d}\lambda_{j}h_{i}(x_{j})\right)-\sum_{i=1}^{2d+1}g\left(\sum_{j=1}^{d}\lambda_{j}\psi_{i}(x_{j})\right)\right|+\left|\sum_{i=1}^{2d+1}g\left(\sum_{j=1}^{d}\lambda_{j}\psi_{i}(x_{j})\right)-\sum_{i=1}^{2d+1}\tilde{\phi}\left(\sum_{j=1}^{d}\lambda_{j}\psi_{i}(x_{j})\right)\right|$
$\leq\sum_{i=1}^{2d+1}\left|g\left(\sum_{j=1}^{d}\lambda_{j}h_{i}(x_{j})\right)-g\left(\sum_{j=1}^{d}\lambda_{j}\psi_{i}(x_{j})\right)\right|+\sum_{i=1}^{2d+1}\left|g\left(\sum_{j=1}^{d}\lambda_{j}\psi_{i}(x_{j})\right)-\tilde{\phi}\left(\sum_{j=1}^{d}\lambda_{j}\psi_{i}(x_{j})\right)\right|$
$<\frac{\epsilon}{2(2d+1)}(2d+1)+\frac{\epsilon}{2(2d+1)}(2d+1)=\epsilon,$
where we applied (3.2) to the first term of the last inequality (since (3.5) holds) and (3.3) to the second term of the last inequality.
Finally, we count the number of intrinsic neurons used in $\phi(\mathbf{x})$. See Fig. 2 for a visualization of the network $\phi$. Note that each $\psi_{i}$ requires $36\times 5+3$ neurons, the linear combination involving $\lambda_{1},\dots,\lambda_{d}$ requires 1 neuron, $\tilde{\phi}$ requires $36\times 5$ neurons, and the final summation requires 1 neuron. In conclusion, the number of intrinsic neurons is $(36\times 5+3)(2d+1)+1+36\times 5+1=366d+365$. The proof is completed. ∎
As mentioned before, the EUAF network in [23, Theorem 1] has width $36d(2d+1)$ and depth $11$, resulting in a total number of neurons equal to $5437(d+1)(2d+1)$. This is because the version of KST used in their proof requires $(2d+1)d$ inner functions and $12d+$ 1 outer functions. Employing the version of KST in Theorem 2.1 not only reduces the total number of intrinsic neurons, but it also simplifies the proof of the existence of an EUAF network with a fixed architecture that can approximate any function in $C([a,b]^{d})$ with arbitrary accuracy. One key thing to keep in mind in the above proof is that the range of the linear combination of point evaluations of an EUAF network (that approximates an inner function) has to be $[0,1]$. This can be done by applying $\min\{\max\{\cdot,0\},1\}$, which can be expressed in terms of the EUAF in (2.2), to the output of each of these EUAF networks. Even though the total number of neurons in Theorem 3.1 is larger than that in [17, Theorem 4], the order of the magnitude is the same, $\mathcal{O}(d)$, and the activation function in our network is explicitly known.
Figure 2. A diagram showing the network $\phi$ in Theorem 3.1. $\mathcal{L}_{\boldsymbol{\lambda}}$ and $\mathcal{L}_{\mathbf{1}}$ respectively represent an inner product with the vector $(\lambda_{1},\dots,\lambda_{d})$ and the vector of ones to obtain a scalar output.
Next, we present a family of continuous functions that requires at least width $d$ (or to put differently, at least $d$ neurons) for it to be approximated with arbitrary accuracy. In the following, we assume that the depth is fixed.
Theorem 3.2.
Let $f\in C([-\tfrac{1}{2},\tfrac{1}{2}]^{d})$ such that $f(\mathbf{0})=0$ (i.e., it vanishes at the origin) and $|f(\mathbf{x})|=|f(x_{1},\dots,x_{d})|\geq c$ for some $c>0$ if $x_{j}=\tfrac{1}{2}$ for some $1\leq j\leq d$ (i.e., $|f(\mathbf{x})|$ is bounded away from zero if at least one of its inputs is equal to $\tfrac{1}{2}$). Then, for any given activation function $\sigma$, the $\sigma$ network with width less than $d$, and fixed depth $L\geq 1$ cannot approximate $f$ with arbitrary accuracy.
Proof.
We proof by contradiction. Assume that for each $\epsilon>0$, there is a $\sigma$ network $\phi$ with width $d-1$ and fixed depth $L\geq 1$ such that
$\left|f(\mathbf{x})-\phi(\mathbf{x})\right|<\epsilon\quad\text{for all }\mathbf{x}\in[-\tfrac{1}{2},\tfrac{1}{2}]^{d},$ (3.6)
where $\phi$ is defined as in (2.1) with $N_{i}=d-1$ for $1\leq i\leq L$. More explicitly,
$\phi(\mathbf{x})=\mathbf{\mathcal{L}}_{L}(\sigma(\mathbf{\mathcal{L}}_{L-1}(\sigma(\dots\mathbf{\mathcal{L}}_{1}(\sigma(\mathbf{W}_{0}\mathbf{x}+\mathbf{b}_{0}))\dots)))),$ (3.7)
where $\mathbf{W}_{0}\in\mathbb{R}^{d-1}\times\mathbb{R}^{d}$ and $\mathbf{b}_{0}\in\mathbb{R}^{d-1}$.
Arbitrarily fix $\epsilon>0$. Consider the homogeneous linear system $\mathbf{W}_{0}\mathbf{x}=\mathbf{0}$. Such a system clearly has infinitely many solutions. Let $\tilde{\mathbf{W}}_{0}$ be the reduced row echelon form of $\mathbf{W}_{0}$. Define $\tilde{\mathbf{x}}:=(\tilde{x}_{1},\dots,\tilde{x}_{d})$ such that $\tilde{\mathbf{W}}_{0}\tilde{\mathbf{x}}=\mathbf{0}$. Then, we can categorize the index of each $\tilde{x}_{i}$ into the following three sets
$\mathscr{I}_{1}$ :=$:=\{1\leq i\leq d:\tilde{x}_{i}=0\},$
$\mathscr{I}_{2}$ :=$:=\{1\leq i\leq d:\text{the $i$-th nonzero row of $\tilde{\mathbf{W}}_{0}$ whose leading entry is }1\},$
$\mathscr{I}_{3}$ :=$:=\{1,\dots,d\}\backslash(\mathscr{I}_{1}\cup\mathscr{I}_{2}).$
Note that the cardinality of the set $\mathscr{I}_{2}$ is equal to the rank of $\tilde{\mathbf{W}}_{0}$ and the cardinality of the set $\mathscr{I}_{1}$ is greater than or equal to zero. Since $\mathbf{W}_{0}\in\mathbb{R}^{d-1}\times\mathbb{R}^{d}$, there is at least one free variable and the number of these free variables is equal to the cardinality of the set $\mathscr{I}_{3}$. Define
$\mathscr{C}:=\left\{c_{ik}\in\mathbb{R}:\tilde{x}_{i}=\sum_{k\in\mathscr{I}_{3}}c_{ik}\tilde{x}_{k},\;i\in\mathscr{I}_{2},\;c_{ik}\neq 0\right\}.$
In fact, each element in $\mathscr{C}$ can be deduced from the entries of $\tilde{\mathbf{W}}_{0}$.
Now, plug $\tilde{\mathbf{x}}$ into (3.6). Assuming that $\tilde{\mathbf{x}}\in[-\tfrac{1}{2},\tfrac{1}{2}]^{d}$, we have
$\epsilon>\left|f(\mathbf{\tilde{\mathbf{x}}})-B\right|\geq\left||f(\mathbf{\tilde{\mathbf{x}}})|-|B|\right|,$ (3.8)
where $B:=\mathbf{\mathcal{L}}_{L}(\sigma(\mathbf{\mathcal{L}}_{L-1}(\sigma(\dots\mathbf{\mathcal{L}}_{1}(\sigma(\mathbf{b}_{0}))\dots))))$ is a constant.
Next, we shall show that indeed we can always pick $\tilde{\mathbf{x}}$ such that $\tilde{\mathbf{W}}_{0}\tilde{\mathbf{x}}=\mathbf{0}$ and $\tilde{\mathbf{x}}\in[-\tfrac{1}{2},\tfrac{1}{2}]^{d}$. Note that we directly have $\tilde{x}_{j}\in[-\tfrac{1}{2},\tfrac{1}{2}]$ for all $j\in\mathscr{I}_{1}$. If $|B|>\epsilon$, then we can set all $\tilde{x}_{j}=0$ for all $j\in\mathscr{I}_{3}$ which subsequently forces all $\tilde{x}_{j}=0$ for all $j\in\mathscr{I}_{2}$. Hence, continuing from (3.8), we have $\epsilon>|B|>\epsilon$ due to the fact that $f(\mathbf{0})=0$, which is a contradiction. Now suppose that $|B|<\epsilon$. Define $\tilde{\mu}:=\operatorname*{arg\,max}_{\mu\in\mathscr{C}}|\mu|$ and set
$\tilde{x}_{\tilde{k}}=\text{sign}(\tilde{\mu})\min\left\{\frac{1}{2},\frac{1}{2}|\tilde{\mu}|^{-1}\right\},$
where $\tilde{k}\in\mathscr{I}_{3}$ is chosen such that $c_{i\tilde{k}}\in\mathscr{C}$ for some $i\in\mathscr{I}_{2}$ and $c_{i\tilde{k}}=\tilde{\mu}$. If there are multiple choices for $\tilde{k}$, then we only pick one of them. Afterwards, we set $\tilde{x}_{j}=0$ for all $j\in\mathscr{I}_{3}$ and $j\neq\tilde{k}$. We can immediately conclude that $|\tilde{x}_{j}|\leq\tfrac{1}{2}$ for all $j\in\mathscr{I}_{2}\cup\mathscr{I}_{3}$. Moreover, this guarantees the existence of at least one $i\in\mathscr{I}_{2}$ such that $\tilde{x}_{i}=\tfrac{1}{2}$. In particular, it corresponds to $i\in\mathscr{I}_{2}$ such that $c_{i\tilde{k}}\in\mathscr{C}$ with $c_{i\tilde{k}}=\tilde{\mu}$. Continuing from (3.8) and using the assumption that $|B|<\epsilon$, we have
$\epsilon>\left||f(\mathbf{\tilde{\mathbf{x}}})|-|B|\right|\geq|f(\mathbf{\tilde{\mathbf{x}}})|-|B|\geq|f(\mathbf{\tilde{\mathbf{x}}})|-\epsilon\geq c-\epsilon,$
where we used our assumption that $|f(\mathbf{x})|=|f(x_{1},\dots,x_{d})|\geq c$ for some $c>0$ if $x_{j}=\tfrac{1}{2}$ for some $1\leq j\leq d$. Therefore, we have a contradiction. The proof is completed. ∎
We provide a concrete example of a function that satisfies the conditions stated in the above theorem.
Example 3.3.
Let $f(\mathbf{x})=\sum_{j=1}^{d}c_{j}h_{j}(x_{j})$, where for all $1\leq j\leq d$, $c_{j}>0$, $x_{j}\in[-\tfrac{1}{2},\tfrac{1}{2}]$, and $h_{j}$ is a nonnegative continuous function such that $h_{j}(0)=0$ and $h_{j}(\tfrac{1}{2})\neq 0$. Clearly, $f\in C([-\tfrac{1}{2},\tfrac{1}{2}]^{d})$, $f(\mathbf{0})=0$, and $|f(\mathbf{x})|=|f(x_{1},\dots,x_{d})|\geq\left(\min_{1\leq j\leq d}c_{j}\right)\left(\min_{1\leq j\leq d}h_{j}(\tfrac{1}{2})\right)$ if $x_{j}=\tfrac{1}{2}$ for some $1\leq j\leq d$. The above theorem states that for any given activation function $\sigma$, the $\sigma$ network with width less than $d$, and fixed depth $L\geq 1$ cannot approximate $f$ with arbitrary accuracy.
Theorem 3.2 presents a family of continuous functions, which cannot be approximated with arbitrary accuracy, when we let the width to be $d-1$ and fix the depth to be $L\geq 1$. This implies that the $\sigma$ network actually requires at least width $d$ and consequently at least $d$ intrinsic neurons to achieve the desired approximation property. Theorems 3.1 and 3.2 combined prove that the requirement of $\mathcal{O}(d)$ fixed intrinsic neurons for approximating functions in $C([a,b]^{d})$ is optimal in the sense that it grows linearly with the input dimension $d$.
Acknowledgment
M. M. was partially supported by NSERC Postdoctoral Fellowship. H. Y. was partially supported by the US National Science Foundation under awards DMS-2244988, DMS-2206333, the Office of Naval Research Award N00014-23-1-2007, and the DARPA D24AP00325-00.